84.)Gone Fishing. Debbie traveled by boat 5 miles upstream to fish in her favorite spot. Because of the 4 - mph current, it took her 20 minutes longer to get there than to return. how fast will her boat go in still water?
Answer by Alan3354(69443)

84.)Gone Fishing. Debbie traveled by boat 5 miles upstream to fish in her favorite spot. Because of the 4 - mph current, it took her 20 minutes longer to get there than to return. how fast will her boat go in still water?
-----------------------------
time = distance/speed
time upstream = 5/(b - 4) b is the boat's speed
time downstream = 5/(b+4) = time upstream - 1/3 (20 mins is 1/3 hours)
---------------
5/(b+4) = 5/(b-4) - 1/3
Multiply by 3*(b+4)*(b-4)
15*(b-4) = 15*(b+4) - b^2 + 16
15b-60 = 15b+ 60 -b^2 + 16
-60 = 76 - b^2
b^2 = 136
b = 11.66 mph